# ImagenHub: Standardizing the evaluation of conditional image generation models

### 03 Dec 2023 v3

@@ 1 Introduction
ImagenHub Inference Library. We built a ImagenHub library111it’s similar to Huggingface libraries (Wolf et al., 2019; von Platen et al., 2022) to evaluate all the conditional image generation models. We ported the highly dispersed codebase from the existing works and then standardized them into a unified format. During inference, we fixed the hyper-parameters and the prompt format to prevent per-instance prompt or hyper-parameter tuning, which makes the inference of different models fair and reproducible. The library is designed to be easily extendable.
@@ 4.1 Human Evaluation Metrics
Table 1: Rating guideline for computing the SC and PQ score. Detail in subsection A.2
@@ 4.2 Dataset and available models
Model #Params Runtime Keywords
@@
Table 3: Overview of all the evaluated models and their parameter size and runtime. The models are listed in chronological order.
@@ 5 Experimental Results
Experiment Setup. All the models either used the default setting from the official implementation or the setting suggested in HuggingFace documentation  (von Platen et al., 2022). We disabled negative prompts and any prompt engineering tricks to ensure a fair comparison. We conducted human evaluation by recruiting participants from Prolific to rate the images, and our own researchers also took part in the image rating process. We assigned 3 raters for each model and computed the SC score, PQ score, and Overall human score. Then we computed the Fleiss kappa  (Fleiss & Cohen, 1973) for each mentioned score. We also computed Krippendorff’s Alpha  (Krippendorff, 2011), which is expected to yield a higher value than Fleiss kappa. This distinction arises from the nature of the rating categories, with Fleiss’ Kappa assuming nominal categories and Krippendorff’s Alpha accommodating ordinal inputs. Both Fleiss’ Kappa and Krippendorff’s Alpha are bounded within the range of [-1, 1], where the value $>0$ indicates an agreement and closer proximity to 1 indicates a higher degree of agreement.
@@ 5.1 Discovery and Insights
Text-Guided Image Generation. We observe that all models are able to generate high-quality images. Regarding semantic consistency, all models have a good understanding of the general prompts, while Stable Diffusion XL is better at understanding complex prompts. For example, it exhibits a high degree of accuracy and detail on the prompt “A panda making latte art.” while other models often misunderstood the prompt as “panda latte art”.
@@ 5.2 Ablation Study
Method of overall human score computation. We set the overall score of the model as the geometric mean of SC and PR score (i.e. $O=\sqrt{SC\times PQ}$). But we also explored the weighted sum setting $O=\alpha\times SC+\beta\times PQ$, where both $\alpha$ and $\beta$ are in [0, 1]. We experimented and found that the weighted sum setting yields a different ranking in the models. Take the text-guided image editing task as an example, Text2Live outperforms CycleDiffusion in the weighted sum setting even though we found that CycleDiffusion performs better in the human examination. We investigated and found that a majority of results in Text2Live simply return the input as output (in that case SC=0 and PQ=1). We tried adjusting the weightings but it still failed to reflect the actual performance of the model. Thus we decided to use the geometric mean setting to penalize the models.
Setting $O=\sqrt{SC\times PQ}$ $O=0.5SC+0.5PQ$ $O=0.7SC+0.3PQ$
$O_{Sum}$ $O_{Avg}$ $O_{Sum}$ $O_{Avg}$ $O_{Sum}$ $O_{Avg}$
MagicBrush 83.51 0.47 103.75 0.58 98.85 0.55
CycleDiffusion 24.89 0.14 65.25 0.36 51.28 0.29
DiffEdit 1.71 0.01 22.08 0.12 14.38 0.08
Text2Live 4.08 0.02 75.25 0.42 46.75 0.26
Table 5: Comparison on overall human score computation setting.
Design choice of human evaluation metric range. When it comes to human evaluation on a massive scale, it’s essential to create a system that’s easy to understand and quick to use. In this investigation, we undertake an exploration into how different settings in this evaluation method can affect the results. Initially, our approach entailed the utilization of a range encompassing [0, 0.5, 1, 2] for both the Semantic Consistency (SC) score and the Perception Quality (PQ) score, where 2 means the image is perfect. However, this configuration yielded suboptimal results in the Fleiss Kappa. Subsequently, an alternative configuration was employed, narrowing the range to [0, 1] for both the SC and PQ scores. This adjustment, while accommodating a binary classification, was observed to yield values that were overly polarized and extreme. To find the right balance between keeping values in a reasonable range and making sure the evaluation method is reliable, we resolved to define a range of [0, 0.5, 1] while providing explicit and unambiguous guidelines.
BLIP-Diffusion 0.13 0.41 0.47 0.09 0.29 0.50 0.05 0.40 0.40
DreamEdit 0.50 0.33 0.52 0.32 0.26 0.55 0.21 0.38 0.38
PhotoSwap 0.62 0.35 0.46 0.36 0.25 0.50 0.29 0.37 0.37
@@ A.1 More metrics results
Table 8: Metrics correlation. we inverted the signs for metric that is the lower the better.
@@ A.7 Visualization Tools (Imagen Museum)
%**** appendix.tex Line 475 ****prompt = dummy_data[’prompt’]
@@
Appendix B Datasheet
B.1 Continuous update of Human evaluation score
Model $SC_{Avg}$ $PQ_{Avg}$ Overall $\text{Fleiss}_{\bar{O}}$ $\text{Kd}_{\bar{O}}$
Text-guided Image Generation
Dalle-3 0.79$\pm$0.02 0.79$\pm$0.14 0.76$\pm$0.08 0.19 0.34
Midjourney 0.67$\pm$0.06 0.92$\pm$0.06 0.73$\pm$0.07 0.34 0.51
DeepFloydIF 0.65$\pm$0.02 0.62$\pm$0.06 0.59$\pm$0.02 0.32 0.51
Stable Diffusion XL 0.62$\pm$0.03 0.64$\pm$0.05 0.59$\pm$0.03 0.37 0.61
Dalle-2 0.58$\pm$0.04 0.62$\pm$0.06 0.54$\pm$0.04 0.27 0.40
OpenJourney 0.53$\pm$0.02 0.59$\pm$0.05 0.50$\pm$0.02 0.30 0.47
Stable Diffusion 2.1 0.56$\pm$0.02 0.53$\pm$0.05 0.50$\pm$0.03 0.38 0.50
Mask-guided Image Editing
SDXL-Inpainting 0.49$\pm$0.05 0.51$\pm$0.02 0.37$\pm$0.05 0.50 0.72
SD-Inpainting 0.28$\pm$0.04 0.27$\pm$0.10 0.17$\pm$0.07 0.31 0.49
GLIDE 0.20$\pm$0.05 0.48$\pm$0.06 0.16$\pm$0.05 0.33 0.56
BlendedDiffusion 0.12$\pm$0.03 0.11$\pm$0.03 0.05$\pm$0.02 0.36 0.44
Text-guided Image Editing
MagicBrush 0.51$\pm$0.01 0.65$\pm$0.06 0.47$\pm$0.02 0.44 0.67
InstructPix2Pix 0.29$\pm$0.01 0.70$\pm$0.06 0.27$\pm$0.02 0.55 0.74
Prompt-to-prompt 0.17$\pm$0.05 0.55$\pm$0.09 0.15$\pm$0.06 0.36 0.53
CycleDiffusion 0.17$\pm$0.03 0.56$\pm$0.11 0.14$\pm$0.04 0.41 0.63
SDEdit 0.04$\pm$0.03 0.56$\pm$0.12 0.04$\pm$0.03 0.13 0.13
Text2Live 0.02$\pm$0.01 0.82$\pm$0.04 0.02$\pm$0.02 0.10 0.17
DiffEdit 0.02$\pm$0.01 0.23$\pm$0.04 0.01$\pm$0.01 0.24 0.24
Pix2PixZero 0.01$\pm$0.00 0.48$\pm$0.09 0.01$\pm$0.01 0.37 0.37
Subject-driven Image Generation
SuTI 0.64$\pm$0.11 0.68$\pm$0.08 0.58$\pm$0.12 0.20 0.39
DreamBooth 0.51$\pm$0.08 0.93$\pm$0.02 0.55$\pm$0.11 0.37 0.60
BLIP-Diffusion 0.29$\pm$0.04 0.93$\pm$0.04 0.35$\pm$0.06 0.22 0.39
TextualInversion 0.21$\pm$0.04 0.74$\pm$0.08 0.21$\pm$0.05 0.35 0.52
DreamBooth-Lora 0.07$\pm$0.01 0.82$\pm$0.07 0.09$\pm$0.01 0.29 0.37
Subject-driven Image Editing
PhotoSwap 0.34$\pm$0.02 0.65$\pm$0.04 0.36$\pm$0.02 0.35 0.46
DreamEdit 0.31$\pm$0.03 0.61$\pm$0.03 0.32$\pm$0.03 0.33 0.52
BLIP-Diffusion 0.09$\pm$0.03 0.70$\pm$0.02 0.09$\pm$0.03 0.41 0.47
Multi-concept Image Composition
CustomDiffusion 0.26$\pm$0.01 0.86$\pm$0.05 0.29$\pm$0.01 0.73 0.88
DreamBooth 0.11$\pm$0.02 0.78$\pm$0.02 0.13$\pm$0.02 0.61 0.71
TextualInversion 0.04$\pm$0.01 0.74$\pm$0.05 0.05$\pm$0.01 0.62 0.77
Control-guided Image Generation
ControlNet 0.42$\pm$0.05 0.19$\pm$0.04 0.23$\pm$0.04 0.37 0.57
UniControl 0.38$\pm$0.07 0.20$\pm$0.06 0.23$\pm$0.07 0.36 0.58
Table 10: Updated table with extra models.
gg==" alt="[LOGO]">

### 10 Mar 2024 v4

@@ 1 Introduction
ImagenHub Inference Library. We built a ImagenHub library111it’s similar to Huggingface libraries (Wolf et al., 2019; von Platen et al., 2022) to evaluate all the conditional image generation models. We ported the highly dispersed codebase from the existing works and then standardized them into a unified format. During inference, we fixed the hyper-parameters and the prompt format to prevent per-instance prompt or hyper-parameter tuning, which makes the inference of different models fair and reproducible. The library is designed to be easily extendable. Our Appendix A.7A.8, and A.9 show how a third party and researchers can benefit from our work.
@@ 4.1 Human Evaluation Metrics
Table 1: Rating guideline for computing the SC and PQ score. Detail in subsection A.2.
@@ 4.2 Dataset and available models
Model #Params Runtime Keywords for technical detail
@@
Dalle-3 (openai, 2023) - - Recaptioned training data
MidJourney (midjourney, 2023) - - -
@@
Table 3: Overview of all the evaluated models and their parameter size and runtime. The models are listed in chronological order. The number of parameters and runtime for Dalle-3 and MidJourney are unknown. MidJourney is not opensource and does not have a whitepaper in technical detail.
@@ 5 Experimental Results
Experiment Setup. All the models either used the default setting from the official implementation or the setting suggested in HuggingFace documentation  (von Platen et al., 2022). We disabled negative prompts and any prompt engineering tricks to ensure a fair comparison. We conducted human evaluation by recruiting participants from Prolific to rate the images, and our own researchers also took part in the image rating process. We assigned 3 raters for each model and computed the SC score, PQ score, and Overall human score. Then we computed the Fleiss kappa  (Fleiss & Cohen, 1973) for each mentioned score. We also computed Krippendorff’s Alpha  (Krippendorff, 2011), which is expected to yield a higher value than Fleiss kappa. This distinction arises from the nature of the rating categories, with Fleiss’ Kappa assuming nominal categories and Krippendorff’s Alpha accommodating ordinal inputs. Both Fleiss’ Kappa and Krippendorff’s Alpha are bounded within the range of [-1, 1], where the value $>0$ indicates an agreement and closer proximity to 1 indicates a higher degree of agreement. The design choices are explained in subsection 5.2 and Appendix A.6.
@@ 5.1 Discovery and Insights
Text-Guided Image Generation. We observe that all models are able to generate high-quality images. Regarding semantic consistency, all models have a good understanding of the general prompts, while Stable Diffusion XL is better at understanding complex prompts. For example, it exhibits a high degree of accuracy and detail on the prompt “A panda making latte art.” while other models often misunderstood the prompt as “panda latte art”. DALLE-3 and Midjourney understand most of the complex prompts than other models, and DALLE-3 is slightly better on conflicting prompts.
@@ 5.2 Ablation Study
Method of overall human score computation. We set the overall score of the model as the geometric mean of SC and PR score (i.e. $O=\sqrt{SC\times PQ}$). But we also explored the weighted sum setting $O=\alpha\times SC+\beta\times PQ$, where both $\alpha$ and $\beta$ are in [0, 1]. We experimented and found that the weighted sum setting yields a different ranking in the models. Take the text-guided image editing task as an example, as in Appendix Table 7, Text2Live outperforms CycleDiffusion in the weighted sum setting even though we found that CycleDiffusion performs better in the human examination. We investigated and found that a majority of results in Text2Live simply return the input as output (in that case SC=0 and PQ=1). We tried adjusting the weightings but it still failed to reflect the actual performance of the model. Thus we decided to use the geometric mean setting to penalize the models.
Design choice of human evaluation metric range. When it comes to human evaluation on a massive scale, it’s essential to create a system that’s easy to understand and quick to use. In this investigation, we undertake an exploration into how different settings in this evaluation method can affect the results, showing in Appendix Table 8. Initially, our approach entailed the utilization of a range encompassing [0, 0.5, 1, 2] for both the Semantic Consistency (SC) score and the Perception Quality (PQ) score, where 2 means the image is perfect. However, this configuration yielded suboptimal results in the Fleiss Kappa. Subsequently, an alternative configuration was employed, narrowing the range to [0, 1] for both the SC and PQ scores. This adjustment, while accommodating a binary classification, was observed to yield values that were overly polarized and extreme. To find the right balance between keeping values in a reasonable range and making sure the evaluation method is reliable, we resolved to define a range of [0, 0.5, 1] while providing explicit and unambiguous guidelines.
@@ A.1 More metrics results
Table 6: Metrics correlation. we inverted the signs for metric that is the lower the better. The computable metrics all hold different assumptions which only contribute part of the aspects in human evaluation. This makes a low correlation with human evaluations.
Setting $O=\sqrt{SC\times PQ}$ $O=0.5SC+0.5PQ$ $O=0.7SC+0.3PQ$
$O_{Sum}$ $O_{Avg}$ $O_{Sum}$ $O_{Avg}$ $O_{Sum}$ $O_{Avg}$
MagicBrush 83.51 0.47 103.75 0.58 98.85 0.55
CycleDiffusion 24.89 0.14 65.25 0.36 51.28 0.29
DiffEdit 1.71 0.01 22.08 0.12 14.38 0.08
Text2Live 4.08 0.02 75.25 0.42 46.75 0.26
Table 7: Comparison on overall human score computation setting.
BLIP-Diffusion 0.13 0.41 0.47 0.09 0.29 0.50 0.05 0.40 0.40
DreamEdit 0.50 0.33 0.52 0.32 0.26 0.55 0.21 0.38 0.38
PhotoSwap 0.62 0.35 0.46 0.36 0.25 0.50 0.29 0.37 0.37
@@ A.7 Visualization Tools (Imagen Museum)
prompt = dummy_data[’prompt’]